Let f(x) = 29 x^2 + 3x + 4 . Determine a formula for g(x) = f(x+15) .

Answer: g(x) = ??

Steps and answer would be nice. Thank you.

Found 2 solutions by josgarithmetic, Edwin McCravy:
Answer by josgarithmetic(39618)

g(x) will still be a function. You are supposed to replace x with x+15 in the function f, and then simplify.

Answer by Edwin McCravy(20056)

f(x) = 29x2 + 3x + 4

Replace every x by (x+15)

f((x+15)) = 29(x+15)2 + 3(x+15) + 4

f(x+15) = 29(x+15)(x+15) + 3x+45 + 4

f(x+15) = 29(x2+30x+225) + 3x+45 + 4

f(x+15) = 29x2+870x+6525 + 3x+45 + 4

f(x+15) = 29x2+873x+6574

g(x) = f(x+15) = 29x2+873x+6574

Edwin